NOAA Office of Satellite and Product Operations

NOAA 14 SEM Subsystem Summary

Component Status:

Component
Description
Status
DPUSpace environment Monitor - Data Processing UnitGreen
MEPEDSpace environment Monitor - Medium Energy Proton DetectorGreen
TEDSpace environment Monitor - Total Energy DetectorYellow

Subsystem History:

Date Time Component Description
10/03/200618:00:00 Restore SEM to standard operational configuration
09/28/200617:05:00 Major spacecraft attitude perturbation resulted in autonomous battery undervoltage safing macro implementation turning off SEM.
12/16/200315:30:00TEDSet TED electron power supply bias voltage to operational level. Recovery ops complete.
12/16/200313:57:00TEDSEM TED pulse height discriminatinator to level-4 (operational value).
12/16/200313:56:00TEDTED electron and proton channel biases to level 1 to continue recovery ops.
12/16/200313:55:00TEDTurn TED high voltage on to continue recovery.
12/16/200313:53:00TEDTurn TED ON to continue SEM recovery .
12/16/200313:53:00TEDTurn TED off for SEM recovery operations
12/16/200309:45:00TEDSEM TED Autonomous Reset.
03/10/200316:04:00TEDTurn TED high voltage on
03/10/200316:04:00TEDTED electron and proton channel biases to known levels
03/10/200316:04:00TEDSets pulse height discriminatinator levels to operational values
03/10/200316:04:00TEDSet electron power supply bias voltage to operational level
03/10/200316:04:00TEDTurn TED On
03/10/200316:04:00TEDTurn TED off
12/14/200222:12:00TEDTED electron and proton channel biases to known levels
12/14/200222:10:00TEDTurn TED high voltage on
12/14/200222:08:00TEDTurn TED On
12/14/200222:06:00TEDTurn TED off
12/14/200210:40:00TEDSEM autonomous reset, recovery ops to Level 4 followed
12/14/200203:20:00TEDSet electron power supply bias voltage to operational level
12/14/200203:18:00TEDSets pulse height discriminatinator levels to operational values
09/01/200200:00:00TEDSEM Autonomous Reset
10/26/200101:50:00TEDPower cycled SEM TED. Set electron and proton biases to level 1, PHD to level 4, MUX to normal mode and set electron channeltron HV to level 0.
10/25/200120:35:00TEDSEM TED bias reset
02/19/200100:00:00TEDSEM testing stepped through PHD levels 1, 3 and 4.
02/08/200119:00:00TEDSEM PHD levels changed from 2 to 4
01/31/200115:30:00TEDSEM PHD increased
01/24/200100:00:00TEDSEM auto shutdown and restart
01/18/200107:00:00TEDSEM auto shutdown and restart @ 11:15
04/15/199900:00:00TEDSEM 300 KeV electron level channels failed, housekeeping data nominal
11/10/199800:00:00TEDSEM TED attempts to configure back to Level 2 unsuccessful, configured back to Level 0.
10/28/199800:00:00TEDPower cycled SEM TED due to high electron and proton activity. Commanded to level 1
01/20/199500:00:00 SEM one of 4 telescopes is not operational, 12% data loss.
12/30/199410:36:00TEDOn Orbit
12/30/199410:36:00DPUOn Orbit
12/30/199410:36:00MEPEDOn Orbit